OpenStudy (anonymous):

Which of the following inequalities matches the graph? graph of an inequality with a solid line through the points (-2, 0) and (0, 3) with shading below the line y < 1 over 3 x + 3 y less than or greater to 3 over 2 x + 3 y greater than or equal to 3 over 2 x + 3 The correct inequality is not listed.

OpenStudy (texaschic101):

first find the slope slope = (y2 - y1) / (x2 - x1) slope = (3 - 0) / (0 - (-2) slope(m) = 3/2 now take either of your points..(0,3) and sub it into y = mx + b... 3 = 3/2(0) + b 3 = b so your equation is : y = 3/2x + 3 a solid line means there is an equal sign in the problem, and shading below the line means it is less then so your inequality is : y <= 3/2x + 3

OpenStudy (anonymous):

K so my answer is C?